# LetsTalk: Latent Diffusion Transformer for Talking Video Synthesis

### v1

@@
Abstract
* Equal contribution. Correspondence to Ruibo Fu: ruibo.fu@nlpr.ia.ac.cn.
Portrait image animation using audio has rapidly advanced, enabling the creation of increasingly realistic and expressive animated faces. The challenges of this multimodality-guided video generation task involve fusing various modalities while ensuring consistency in timing and portrait. We further seek to produce vivid talking heads. To address these challenges, we present LetsTalk (LatEnt Diffusion TranSformer for Talking Video Synthesis), a diffusion transformer that incorporates modular temporal and spatial attention mechanisms to merge multimodality and enhance spatial-temporal consistency. To handle multimodal conditions, we first summarize three fusion schemes, ranging from shallow to deep fusion compactness, and thoroughly explore their impact and applicability. Then we propose a suitable solution according to the modality differences of image, audio, and video generation. For portrait, we utilize a deep fusion scheme (Symbiotic Fusion) to ensure portrait consistency. For audio, we implement a shallow fusion scheme (Direct Fusion) to achieve audio-animation alignment while preserving diversity. Our extensive experiments demonstrate that our approach generates temporally coherent and realistic videos with enhanced diversity and liveliness. The project page is available at Let’s Talk.
{strip}
[Uncaptioned image][Uncaptioned image]
@@ Abstract
Figure 1: We propose LetsTalk, a diffusion-based transformer for audio-driven portrait image animation. Left: Given a single reference image and audio, LetsTalks can produce a realistic and vivid video aligned with the input audio. Note that each column corresponds to the same audio. The results show that LetsTalk can drive consistent and reasonable mouth motions for the input audio. Right: Generation quality vs. inference time on the HDTF dataset, the circle area reflects the number of parameters of the method. Our LetsTalk achieves the best quality while also being highly efficient in inference, compared to current mainstream diffusion-based methods such as Hallo and AniPortrait. In addition, our base version (LetsTalk-B) achieves performance similar to Hallo with only 8 $\times$ fewer parameters.
1 Introduction
Animating a portrait image with speech, also known as talking head generation, is a challenging task that aims to generate a speaking portrait from a single image with audio. Naturally driving the reference portrait according to the given audio shows great potential in video gaming, film, digital marketing, etc. However, producing reasonable animation faces the following main challenges: consistency in timing and portrait, audio-animation alignment, and diversity.
To overcome these challenges, several methods [52, 45] adopt a 3D morphable model (3DMM) [4, 26, 3] to represent the reference portrait and use audio signals to drive the 3DMM. They feed the 3D signals (e.g. coefficients, poses) from the driven 3DMM into a neural renderer to generate a talking head. While achieving great consistency in timing and portrait, they tend to produce dull results, such as unblinking eyes and fixed hair, which damage the diversity of animations. In contrast to these 3D-aware methods, some other methods [7, 32, 55, 10, 54] mainly focus on capturing the correspondence between the given audio and related 2D movements, like facial muscle and lip motions, directly aligning animation with audio. They implicitly learn the correspondence between audio and animation in the latent space, animating the portrait image by feeding forward the audio. However, these methods still struggle to disentangle the components representing the static and dynamic aspects of the face, even hindering the temporal consistency and diversity of videos.
Recently, diffusion models [18, 13, 34] have emerged as a promising technique to generate realistic 2D images [34, 51] and have even achieved remarkable results in the video generation [29, 28]. Notably, AnimateAnyone [21] incorporates motion modules and ReferenceNet into diffusion models to generate realistic and pose-controllable human animation, which verifies that the end-to-end diffusion models can seamlessly capture the correspondence between the condition signal and human animation. Inspired by it, recent methods [47, 9, 11] replace the pose signals in AnimateAnyone with audio signals and achieve talking head generation. Leveraging diffusion models, these methods excel at generating various unconscious actions, thereby enhancing diversity. While achieving promising results, these methods adopt U-Net as the backbone and lack the flexibility for better fusing multimodality. In addition, although the U-Net backbone works well for image generation, it introduces considerable payloads and struggles with scalability for video generation. Therefore, it is necessary to propose a transformer framework to solve this task and reasonably fuse multimodality, ensuring better portrait consistency and producing more vivid characters.
In this work, we present LetsTalk, a latent diffusion transformer to solve this video generation-like task via the diffusion paradigm. In LetsTalk, we first employ a variational autoencoder (VAE) to reduce this task into a generation task within low-resolution latent space. To enhance the spatial-temporal consistency, we equip each transformer block with temporal and spatial attention modules to capture intra-frame and cross-frame coherence. Regarding multimodal guidance, we define it in this task as the fusion of video generation with reference image and audio sequences. To better exploit multimodality to guide video generation, we summarize three schemes from shallow to deep according to the compactness of the fusion: 1) Direct Fusion, which uses a fixed modality to guide all features in the backbone through a cross-attention mechanism; 2) Siamese Fusion, like ReferenceNet, which feeds modality into a backbone-like network and uses the corresponding features to guide the features in the backbone; 3) Symbiotic Fusion, which concatenates modality with the input at the beginning and achieves guidance only through backbone’s inherent self-attention mechanisms. After summarizing these fusion schemes, we explore the impact of adopting different schemes for the given reference image and audio sequence. To better guide video generation, we use Symbiotic Fusion for the reference image whose modality is close to the video while strongly ensuring the portrait consistency of the generated video. Given the audio, we ask the animation to be aligned with it, generating plausible animation, especially lip motions, and further pursuing diversity (e.g. blinking, head moving, and expressions). Considering the audio modality differs from video, we adopt the Direct Fusion that ensures the audio-animation alignments and avoids ambiguity caused by significant modality differences in the fusion. Additionally, we find that unconscious portrait actions, closely related to diversity, are primarily generated during the denoising of diffusion models, and the shallow fusion scheme effectively preserves the diversity. The empirical results indicate that our innovative framework realizes realistic and high-fidelity talking-head synthesis, advancing the frontiers of realism and expressiveness in audio-driven video generation. To the best of our knowledge, our method is the first work to introduce the diffusion transformer for audio-driven portrait animation.
@@ 1 Introduction
  • We present LetsTalk that pioneers the utilization of transformers in diffusion-based, audio-driven portrait animation by introducing modular spatial-temporal attention mechanisms;
@@
  • We summarize and deeply analyze various multimodal fusion schemes, and employ Symbiotic Fusion and Direct Fusion to fuse the input reference portrait and audio respectively, reasonably guiding video generation via multimodality;
  • Our experiments on standard benchmarks demonstrate the superiority of our methods to other approaches in driving portrait animation via audio.
2 Related Work
2.1 Diffusion Models for Video Generation
@@
Diffusion models have demonstrated exceptional potential across various domains, particularly in image [18, 13] and video generation [17, 21, 20, 57]. A notable example is Stable Diffusion [34], which utilizes a UNet architecture to iteratively produce high-quality images from text prompts after extensive training on large-scale text-image datasets [37]. This approach has been extended to more complex tasks like video generation [17, 21], where models such as Video Diffusion Models (VDM)[20] apply the diffusion process with space-time factorized training, and ImagenVideo[19] enhances VDM with cascaded diffusion models for high-resolution outputs. Further advances, including VideoComposer [43], AnimateDiff [17], and VideoCrafter [6], focus on synthesizing videos by conditioning the diffusion process on images and blending image latents. A significant direction in the evolution of diffusion models is the incorporation of patch sequences of images and transformers, exemplified by the Diffusion Transformer (DiT) architecture [31], which facilitates scaling to larger datasets and model parameters. Recent works, such as Latte [29] and VDT [28], have extended the DiT architecture for video generation by incorporating temporal mechanisms, achieving strong results by training diffusion models from scratch. Building on this, our study also adopts an end-to-end approach for high-fidelity, temporally consistent video generation.
2.2 Audio-driven Portrait Image Animation
@@
In recent years, the field of portrait image animation has advanced significantly, especially in generating realistic and expressive talking head animations from static images paired with audio inputs. Early efforts primarily focused on improving lip synchronization, with models like LipSyncExpert [32] achieving high accuracy for specific identities. Subsequently, methods such as ATVGnet [7] and MakeItTalk [55] introduced facial landmarks and disentangled representations to improve video frame generation. The integration of 3D information in models like SadTalker [52], Genface [50], and DiffTalk [38] further improved head movements and expressions. A significant advancement came with the introduction of diffusion models, including Diffused Heads [39], DreamTalk [30], and VividTalk [40], which pushed the field forward by producing high-fidelity, natural talking head videos across diverse identities, focusing on improved lip synchronization and varied speaking styles. Recent innovations like VASA-1 [48], EMO [41] and AniPortrait [45] have focused on creating emotionally expressive talking faces, capturing a wide range of facial nuances and movements. AniTalker [27] introduced a universal motion model for realistic facial movements with minimal labeled data, enabling dynamic avatars. LivePortrait [16] extends the implicit-keypoint framework, balancing efficiency and controllability. Most recently, Hallo [47, 11] and X-Portrait [46] have adopted the latent diffusion framework pioneered in AnimateAnyone [21] or ControlNet [51], adding more control over elements such as lip movements, expressions, and poses.
2.3 Conditionally Guided Diffusion Model
@@
The impressive generative capabilities of diffusion models have made the integration of control information during generation a crucial area of research. Text-guided control methods focus on refining prompts, manipulating CLIP features, and modifying cross-attention mechanisms [33, 1, 5, 49, 24]. ControlNet [51] enhances large pretrained text-to-image diffusion models by incorporating spatially localized, task-specific conditioning. In video generation, AnimateAnyone [21] introduces the ReferenceNet to ensure the consistency of fine-grained appearance details from reference images. Gen-1 [15] integrates structural guidance through depth maps and cross-attention, while MCDiff [8], LaMD [22], and MotionCtrl [44] focus on motion-guided video generation, enhancing the realism of human actions. For talking video synthesis, previous approaches primarily relied on image diffusion models to generate dynamic portraits frame by frame, starting from reference portraits [38, 39]. Recent methods often incorporate ReferenceNet for processing reference portraits and cross-attention mechanisms for handling audio information [41, 47]. While control over both images and videos in latent diffusion models has seen significant advancements, diffusion transformers remain underdeveloped in this area, particularly when it comes to handling multiple modalities (e.g. image, video, audio). To address this gap, we explore various diffusion conditioning strategies tailored to the transformer architecture, aiming to achieve better control and enhanced generation quality.
(a) Overview of LetsTalk
(b) Transformer Block
Figure 2: The overview of our method (a) and the illustration of our designed transformer block (b). For better illustration, we omit the timestep encoder and Layer Norm in (b). LetsTalk integrates transformer blocks equipped with both temporal and spatial attention modules, designed to capture intra-frame spatial details and establish temporal correspondence across time steps. After obtaining portrait and audio embeddings, Symbiotic Fusion is used for fusing the portrait embedding and Direct Fusion is for fusing the audio embedding. Notably, we repeat the portrait embedding along the frame axis to make it have the same shape as the noise embedding.
3 Method
We present LetsTalk, a diffusion transformer for portrait image animation driven by speech, as shown in Fig. 2. In this section, we first briefly introduce diffusion models and the formulation of our task in Sec. 3.1. Then, we delve into the details of our transformer backbone in Sec. 3.2. Finally, we summarize multimodal fusion schemes and present our adopted strategy for fusing the input reference portrait and audio in Sec. 3.3.
3.1 Preliminaries
Diffusion Formulation The diffusion process incrementally corrupts data $x_{0}$ with Gaussian noise: $q(x_{t}|x_{0})=\mathcal{N}(x_{t};\sqrt{\bar{\alpha}_{t}}x_{0},(1-\bar{\alpha}_{t})\mathbf{I})$, where $\bar{\alpha}_{t}$ are predetermined hyperparameters. By reparameterizing, we obtain: $x_{t}=\sqrt{\bar{\alpha}_{t}}x_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon_{t}$, where $\epsilon\sim\mathcal{N}(0,1)$, forming a Markov chain over $T$ steps.
$\begin{split}\vspace{-0.3cm}\mathcal{L}(\theta)=&\sum_{t}\mathcal{D}_{\text{KL}}(q^{*}(x_{t-1}|x_{t},x_{0})\,\|\,p_{\theta}(x_{t-1}|x_{t}))\\ &-p(x_{0}|x_{1}).\end{split}$ (1)
$\mathcal{L}_{\text{simple}}(\theta)=\|\epsilon_{\theta}(x)-\epsilon(x)\|_{2}^{2}.$ (2)
Diffusion Transformer To better handle the multi-modality fusion in the talking head task, we are motivated to extend the Diffusion Transformer (DiT) in our framework. DiT builds upon the Vision Transformer (ViT) architecture, which processes sequences of image patches. This sequential patch-based processing mechanism natively supports the fusion of multi-modality sufficiently and efficiently. Different from ViT, DiT incorporates a variational autoencoder (VAE) [25] in the framework, thus runs in the latent space, rather than pixel space. Such design enables DiT to handle the expensive high-resolution generation task, which brings more obvious benefits in the video generation task.
Since we follow DiT and introduce a pretrained VAE in our framework, our goal is reduced from generating a video $\bm{V}\in\mathbb{R}^{F\times\hat{H}\times\hat{W}\times 3}$ to generating a latent feature $\bm{F}\in\mathbb{R}^{F\times H\times W\times C}$, where $F$ denotes the number of video frames333 The encoder in VAE downsamples the image by a factor of $8$, and $\hat{H}=8\times H,\hat{W}=8\times W$ We first separate the latent feature $\bm{F}$ into non-overlapping square patches along the spatial dimensions (i.e., $H$ and $W$) and obtain a patch sequence of length $F\times P$. For each patch, we flatten it and use an MLP to project it into the latent space as $\bm{Z}\in\mathbb{R}^{F\times P\times d}$, where $d$ is the dimension of the latent space. Then we add spatial and temporal positional embeddings [42] to the latent feature and make it sensitive to spatial and temporal information. Finally, we feed the processed latent feature into transformers and perform denoising.
(a) Direct Fusion
(b) Siamese Fusion
(c) Symbiotic Fusion
Figure 3: Illustration of three multimodal fusion schemes, our transformer backbone is formed by the left-side blocks. (a) Direct Fusion. Directly feeding condition into each block’s cross-attention module; (b) Siamese Fusion. Maintaining a similar transformer and feeding the condition into it, extracting the corresponding features to guide the features in the backbone; (c) Symbiotic Fusion. Concatenating modality with the input at the beginning, then feeding it into the backbone, achieving fusion via the inherent self-attention mechanisms.
3.2 Efficient Spatial-Temporal-Aware Transformer Backbone
Given a sequence of latent features $\bm{Z}\in\mathbb{R}^{F\times P\times d}$, we feed it into a transformer and predict the noise to be removed. For video generation, we extend the backbone in DiT that only consists of spatial transformer blocks. It is straightforward to flatten the $\bm{Z}$ into $F\times P$ tokens with shape $(F\times P,d)$ and repeatedly conduct a self-attention mechanism to achieve spatiotemporal fusion. However, this approach results in a huge payload for computing the attention matrix with shape $(F\times P,F\times P)$, as it grows with the square of the increase in resolution or frames.
To address this problem, we propose to achieve spatiotemporal fusion in an “interleaved fusion” manner, implemented by alternately stacking spatial and temporal blocks. Considering that $F$ refers to the temporal dimension (i.e. the number of frames) and $P$ refers to the spatial dimension (i.e. the number of patches in each frame), to feed the feature into the spatial block, we maintain the feature with shape $(F,P,d)$ and calculate the attention matrix with shape $(F,P,P)$ to capture spatial information exclusively among patches in the same frame. Similarly, to feed the feature into the temporal block, we maintain the feature with shape $(P,F,d)$ and calculate the attention matrix with shape $(P,F,F)$ to capture temporal information among frames in the same patch. Compared with the straightforward scheme that produces an attention matrix with $F^{2}\times P^{2}$ parameters, our scheme significantly reduces the amount to $F\times P^{2}+P\times F^{2}$. Our experiments verify the efficacy of our backbone extension in terms of image quality and temporal consistency.
3.3 Multimodal Fusion Schemes for Portrait and Speech Audio

### v2

# Efficient Long-duration Talking Video Synthesis with Linear Diffusion Transformer under Multimodal Guidance

@@
Abstract.
Portrait image animation using audio has rapidly advanced, but challenges remain in efficiently fusing multimodal inputs while ensuring temporal and portrait consistency with minimal computational cost. To address this, we present LetsTalk, a LinEar diffusion TranSformer for Talking video synthesis. LetsTalk incorporates a deep compression autoencoder to obtain efficient latent representations, and a spatio-temporal-aware transformer with efficient linear attention to effectively fuse multimodal information and enhance spatio-temporal consistency. We systematically explore and summarize three fusion schemes, ranging from shallow to deep fusion. We thoroughly analyze their characteristics, applicability, and trade-offs, thereby bridging critical gaps in multimodal conditional guidance. Based on modality differences of image, audio, and video generation, we adopt deep (Symbiotic Fusion) for portrait to ensure consistency, and shallow (Direct Fusion) for audio to align animation with speech while preserving motion diversity. To maintain temporal consistency in long-duration video generation, we propose a memory bank mechanism that preserves inter-clip dependencies, effectively preventing degradation across extended sequences. Furthermore, we develop a noise-regularized training strategy that explicitly compensates for DDPM sampling artifacts, significantly improving the model’s robustness in continuous generation scenarios. Our extensive experiments demonstrate that our approach achieves state-of-the-art generation quality, producing temporally coherent and realistic videos with enhanced diversity and liveliness, while maintaining remarkable efficiency through its optimized model design with 8$\times$ fewer parameters.
Talking Head Generation, Multimodal Guidance, Long-duration Generation, Linear Diffusion Transformer, Deep Compression Autoencoder
ccs: Computing methodologies Procedural animationccs: Computing methodologies Computer vision tasksccs: Computing methodologies Neural networks
@@ Abstract.
Figure 1. We introduce LetsTalk, a diffusion-based transformer for audio-driven portrait animation. Given a reference image and audio, LetsTalk generates realistic videos with synchronized mouth motions. As shown in the Left figure, each column corresponds to the same audio, demonstrating consistent and accurate lip movements. The Right figure compares generation quality and inference time on the HDTF dataset, where circle size represents model parameters. LetsTalk achieves superior quality and efficiency, outperforming methods like Hallo and AniPortrait. Notably, our base version (LetsTalk-B) matches Hallo’s performance with 8$\times$ fewer parameters.
[…]
Enjoying the baseball game from the third-base seats. Ichiro Suzuki preparing to bat.
1. Introduction
Animating a portrait image with speech, also known as talking head generation, is a challenging task that aims to generate a speaking portrait from a single image with audio. Naturally driving the reference portrait according to the given audio shows great potential in video gaming, film, digital marketing, etc.. Previous works (Zhang et al., 2023a; Wei et al., 2024) use a 3D morphable model (3DMM) (Blanz and Vetter, 1999; Li et al., 2017; Bao et al., 2021) to represent reference portrait and drive them with audio signals. These approaches input driven 3DMM parameters (e.g., coefficients, poses) into neural renderers, achieving temporal and portrait consistency yet lacking output diversity, often showing static features like unblinking eyes and rigid hair. In contrast, audio-to-2D facial motion methods (Chen et al., 2019; Prajwal et al., 2020; Zhou et al., 2020; Cheng et al., 2022; Zhou et al., 2021) directly align audio with facial dynamics via latent space learning. However, their limited disentanglement of static and dynamic components compromises both temporal consistency and animation diversity.
Recently, diffusion models (Ho et al., 2020; Dhariwal and Nichol, 2021; Rombach et al., 2022) have emerged as a powerful paradigm for generating realistic 2D images (Rombach et al., 2022; Zhang et al., 2023b) and have demonstrated remarkable success in video generation (Ma et al., 2024; Lu et al., 2023). Notably, AnimateAnyone (Hu, 2024) introduces motion modules and ReferenceNet into diffusion models to produce realistic and pose-controllable human animations, showcasing the capability of end-to-end diffusion models to effectively capture the correspondence between conditional signals and human motion. Building on this, recent methods (Xu et al., 2024b; Chen et al., 2024; Cui et al., 2024) have extended this approach by replacing pose signals with audio signals for talking head generation. These methods leverage diffusion models to generate diverse unconscious actions, significantly enhancing the richness of generated content.
Despite these advancements, several critical challenges remain unaddressed. First, current approaches suffer from suboptimal efficiency in audio-driven animation, primarily due to: (1) reliance on auxiliary spatial signals (e.g., predefined movement regions) as a crutch for limited audio controllability, which sacrifices motion diversity and naturalness; (2) redundant parameter overhead from fine-tuning generic text-to-image/video models, despite their misalignment with animation-specific requirements; and (3) inefficient architectures, exemplified by low-compression-rate autoencoders and vanilla attention mechanisms with high computational complexity. Second, despite the focus on conditional video generation, there is a lack of systematic exploration into the intrinsic modeling principles for effectively integrating multimodal conditions, such as reference portraits and driving audio. Lastly, while existing methods attempt to achieve long-duration portrait animation through sequential frame synthesis, they fail to address the progressive accumulation of residual noise in generated frames. This phenomenon inevitably introduces cumulative artifacts (e.g., facial distortions or flickering) that amplify over extended sequences, ultimately degrading output fidelity.
In this work, we introduce LetsTalk, a linear diffusion transformer for efficient generation of realistic, long-duration talking videos under multimodal conditions. Built upon a diffusion transformer as the video generation backbone, LetsTalk replaces conventional text-to-image architectures with an efficient design tailored for audio-driven animation. Key to our approach is a deep compression autoencoder that compresses input frames 32×, drastically reducing the number of latent tokens while preserving critical visual details. Each transformer block incorporates dedicated temporal and spatial modules, where linear attention (Xie et al., 2024a) ensures efficient modeling of intra-frame and cross-frame coherence without sacrificing computational efficiency.
Regarding multimodal guidance, we define it in this task as the fusion of video generation with reference portrait and driving audio. To better exploit multimodality to guide video generation, we first summarize three schemes from shallow to deep according to the compactness of the fusion: 1) Direct Fusion, 2) Siamese Fusion, and 3) Symbiotic Fusion. We analyze these fusion schemes in terms of integration, conditional state, and parameter sharing, systematically summarizing their adaptability to different modalities, and further investigate the underlying factors influencing such adaptability—thereby bridging theoretical gaps in conditional guidance modeling. To better guide video generation effectively, we use Symbiotic Fusion for the reference image, ensuring strong portrait consistency due to its visual similarity to the video. For audio, we employ Direct Fusion to align animations (e.g., lip motion, expressions) with the audio while maintaining diversity in movements (e.g., blinking, head motion).
To enhance long-duration video generation, we incorporate a memory bank mechanism to maintain contextual continuity across sequential video sequences. Furthermore, we propose a novel noise-regularized training strategy and a patch drop augmentation to align the memory bank between training and inference. By explicitly simulating the residual noise accumulation that occurs during inference (where errors propagate as the video extends), our approach learns to compensate for such artifacts during training for robust long-sequence generation.
@@ 1. Introduction
  • We present LetsTalk, a linear diffusion transformer for audio-driven portrait animation, which employs efficient linear attention for spatiotemporal modeling and integrates a deep compression autoencoder to achieve 32 $\times$ latent space reduction while preserving visual fidelity;
@@
  • We systematically analyze three multimodal fusion schemes, characterize their adaptation properties, and employ Symbiotic Fusion and Direct Fusion to fuse the input reference portrait and audio respectively, reasonably guiding video generation via multimodality guidance;
  • We enable long-duration video generation via a memory bank framework with noise-regularized training, aligning memory states between training and inference to mitigate cumulative artifacts in extended sequences;
  • Our experiments demonstrate our method’s superiority in generating realistic, temporally coherent videos with enhanced diversity, achieving an optimal balance between computational efficiency and output quality.
2. Related Work
2.1. Diffusion Models for Video Generation
@@
Diffusion models have demonstrated exceptional potential across various domains, particularly in image (Ho et al., 2020; Dhariwal and Nichol, 2021) and video generation (Guo et al., 2023; Hu, 2024; Ho et al., 2022b; Zhu et al., 2024). A notable example is Stable Diffusion (Rombach et al., 2022), which utilizes a UNet architecture to iteratively produce high-quality images from text prompts after extensive training on large-scale text-image datasets (Schuhmann et al., 2022). This approach has been extended to more complex tasks like video generation (Guo et al., 2023; Hu, 2024), where models such as Video Diffusion Models (VDM)(Ho et al., 2022b) apply the diffusion process with space-time factorized training, and ImagenVideo(Ho et al., 2022a) enhances VDM with cascaded diffusion models for high-resolution outputs. Further advances, including VideoComposer (Wang et al., 2024b), AnimateDiff (Guo et al., 2023), and VideoCrafter (Chen et al., 2023b), focus on synthesizing videos by conditioning the diffusion process on images and blending image latents. A significant direction in the evolution of diffusion models is the incorporation of patch sequences of images and transformers, exemplified by the Diffusion Transformer (DiT) architecture (Peebles and Xie, 2023), which facilitates scaling to larger datasets and model parameters. Recent works, such as Latte (Ma et al., 2024) and VDT (Lu et al., 2023), have extended the DiT architecture for video generation by incorporating temporal mechanisms, achieving strong results by training diffusion models from scratch. Building on this, our study also adopts an end-to-end approach for high-fidelity, temporally consistent video generation.
2.2. Audio-driven Portrait Image Animation
@@
In recent years, the field of portrait image animation has advanced significantly, especially in generating realistic and expressive talking head animations from static images paired with audio inputs. Early efforts primarily focused on improving lip synchronization, with models like LipSyncExpert (Prajwal et al., 2020) achieving high accuracy for specific identities. Subsequently, methods such as ATVGnet (Chen et al., 2019) and MakeItTalk (Zhou et al., 2020) introduced facial landmarks and disentangled representations to improve video frame generation. The integration of 3D information in models like SadTalker (Zhang et al., 2023a), Genface (Ye et al., 2023), and DiffTalk (Shen et al., 2023) further improved head movements and expressions. A significant advancement came with the introduction of diffusion models, including Diffused Heads (Stypułkowski et al., 2024), DreamTalk (Ma et al., 2023), and VividTalk (Sun et al., 2023), which pushed the field forward by producing high-fidelity, natural talking head videos across diverse identities, focusing on improved lip synchronization and varied speaking styles. Recent innovations like VASA-1 (Xu et al., 2024a), EMO (Tian et al., 2024) and AniPortrait (Wei et al., 2024) have focused on creating emotionally expressive talking faces, capturing a wide range of facial nuances and movements. AniTalker (Liu et al., 2024) introduced a universal motion model for realistic facial movements with minimal labeled data, enabling dynamic avatars. LivePortrait (Guo et al., 2024) extends the implicit-keypoint framework, balancing efficiency and controllability. Most recently, Hallo (Xu et al., 2024b; Cui et al., 2024), EchoMimic (Chen et al., 2024) and X-Portrait (Xie et al., 2024b) have adopted the latent diffusion framework pioneered in AnimateAnyone (Hu, 2024), adding more control over elements such as lip movements, expressions, and poses.
2.3. Conditionally Guided Diffusion Model
@@
The impressive generative capabilities of diffusion models have made the integration of control information during generation a crucial area of research. Text-guided control methods focus on refining prompts, manipulating CLIP features, and modifying cross-attention mechanisms (Ramesh et al., 2022; Avrahami et al., 2022; Brooks et al., 2023; Yang et al., 2024; Khachatryan et al., 2023). ControlNet (Zhang et al., 2023b) enhances large pretrained text-to-image diffusion models by incorporating spatially localized, task-specific conditioning. In video generation, AnimateAnyone (Hu, 2024) introduces the ReferenceNet to ensure the consistency of fine-grained appearance details from reference images. Gen-1 (Esser et al., 2023) integrates structural guidance through depth maps and cross-attention, while MCDiff (Chen et al., 2023a), LaMD (Hu et al., 2023), and MotionCtrl (Wang et al., 2024a) focus on motion-guided video generation, enhancing the realism of human actions. For talking video synthesis, previous approaches primarily relied on image diffusion models to generate dynamic portraits frame by frame, starting from reference portraits (Shen et al., 2023; Stypułkowski et al., 2024). Recent methods often incorporate ReferenceNet for processing reference portraits and cross-attention mechanisms for handling audio information (Tian et al., 2024; Xu et al., 2024b). While control over both images and videos in latent diffusion models has seen significant advancements, the theoretical understanding of various conditioning strategies remains limited, particularly in handling multiple modalities (e.g., image, video, audio). To address this gap, we investigate conditioning strategies in diffusion transformers, aiming to improve control and generation quality across modalities.
Figure 2. The illustration of our method. LetsTalk integrates a deep compression autoencoder that efficiently reduces spatial redundancy while preserving temporal features, and transformer blocks equipped with dual temporal and spatial attention modules, enabling the capture of intra-frame spatial details while establishing temporal correspondences across frames. After obtaining portrait and audio embeddings, Symbiotic Fusion is used for fusing the portrait embedding and Direct Fusion is for fusing the audio embedding. Notably, we repeat the portrait embedding along the frame axis to make it have the same shape as the noise embedding. To support long-duration video generation, we introduce a memory bank module to preserve temporal consistency across video sequences. Additionally, we devise a novel training strategy to to align memory bank between training and inference. By explicitly simulating the residual noise during inference, our approach learns to compensate for such artifacts, ensuring robust long-duration generation.
3. Method
We present LetsTalk, a diffusion transformer for audio-driven portrait animation, as illustrated in Fig. 2. In this section, we first detail our efficient transformer architecture in Sec. 3.1. Next, we outline multimodal fusion schemes in Sec. 3.2 and describe our strategy for integrating reference portraits with audio in Sec. 3.3. Finally, we introduce our noise-regularized training approach for long-duration video generation Sec. 3.4.
3.1. Efficient Spatial-Temporal Transformer with Compression Autoencoder
Deep Compression Autoencoder: Building upon latent diffusion models, we employ a deep compression autoencoder (Xie et al., 2024a) that downsamples images by a factor of 32 (vs. 8 in conventional VAEs), while maintaining reconstruction quality comparable to SDXL’s autoencoder through advanced architectural designs. This design is motivated by two key insights: (1) high-resolution video frames contain substantial redundant information that can be safely compressed, and (2) for audio-driven generation, separating compression from denoising enables specialized processing - the autoencoder preserves crucial temporal features (e.g., lip-sync dynamics) while the diffusion model focuses on audio-visual alignment.
The compressed latent feature $\mathcal{F}\in\mathbb{R}^{F\times H\times W\times C}$ is processed through non-overlapping patches, where spatial-temporal positional embeddings (Vaswani, 2017) preserve structural relationships. This compact representation enables efficient transformer processing while reducing memory requirements by 4$\times$ compared to conventional approaches.
Interleaved Fusion with Linear Attention: Given a sequence of latent features $\bm{Z}\in\mathbb{R}^{F\times P\times d}$, we employ a transformer architecture to predict the noise for removal. While the original DiT (Peebles and Xie, 2023) processes individual frames independently, our video generation framework extends this through interleaved spatiotemporal fusion. The naive approach of flattening $\bm{Z}$ into $(F\times P,d)$ tokens with standard attention incurs prohibitive quadratic complexity $\mathcal{O}(F^{2}P^{2})$ for high-resolution video generation.
Our architecture employs a dual-block decomposition strategy that alternates between spatial and temporal modeling, significantly reducing computational complexity from $\mathcal{O}(F^{2}P^{2})$ to $\mathcal{O}(F^{2}+P^{2})$. To reduce computational complexity, we replace the traditional softmax attention with ReLU linear attention (Katharopoulos et al., 2020). For spatial modeling, the input is first reshaped into $(F,P,d)$ dimensions to compute intra-frame attention using:
(1) $\bm{Z_{s}}=\text{LinearAttn}(\bm{Q}_{s},\bm{K}_{s},\bm{V}_{s})=\frac{\sum_{j=1}^{N}\phi(\bm{V}_{j})\phi(\bm{K}_{j})^{\top}\bm{Q}_{s}}{\sum_{j=1}^{N}\phi(\bm{K}_{j})^{\top}\bm{Q}_{s}}$
where $\bm{Q}_{s},\bm{K}_{s},\bm{V}_{s}\in\mathbb{R}^{F\times P\times d}$, $\phi(x)=\text{ReLU}(x)$. The linear attention mechanism further optimizes this by reusing shared terms across all queries: $\sum_{j=1}^{N}\text{ReLU}(\bm{K}_{j})\bm{V}_{j}^{T}\in\mathbb{R}^{d\times d}$.
For temporal modeling, the tensor to $(P,F,d)$ dimensions. Since $F\ll P$, we preserve the expressive power of standard softmax attention for inter-frame interactions: $\bm{Z_{t}}=\text{Attn}(\bm{Q}_{t},\bm{K}_{t},\bm{V}_{t})=\frac{\bm{V}_{t}(\bm{K}_{t}^{\top}\bm{Q}_{t})}{\sqrt{d}}$. Through this hybrid decomposition with linear attention, the overall complexity is reduced to $\mathcal{O}(F^{2}+Pd))$ achieving near-linear efficiency while preserving global receptive fields.
Mix-FFN block In spatial modeling, linear attention converges slower despite matching final performance. Following SANA (Xie et al., 2024a), we improve training efficiency by replacing the original MLP-FFN with Mix-FFN, which combines an inverted residual block, 3×3 depth-wise convolution, and GLU. The depth-wise convolution enhances local feature capture, compensating for ReLU linear attention’s weaker locality. Our experiments demonstrate this architecture achieves superior computational efficiency without compromising visual quality or temporal consistency compared to quadratic-complexity baselines in Fig. 1.
Figure 3. Multimodal fusion schemes: (a) Direct Fusion injects conditions via cross-attention modules; (b) Siamese Fusion uses parallel transformer for feature guidance; (c) Symbiotic Fusion achieves fusion through input concatenation and self-attention. The backbone architecture (left-side blocks) remains consistent across all approaches.
Schemes Integration State Params Modality Adapt.
Direct Cross-attention Static Dedicated Weak alignment
Siamese Self-attention/Add Dynamic Dedicated Flexible
Symbiotic Self-attention Dynamic Shared Strong
Table 1. Comparative analysis of three fusion schemes across three key dimensions: integration methods, condition states, and parameters sharing implementation
3.2. Analysis of Multimodal Fusion Schemes